### v1

# Cobham’s Theorem for the Gaussian integers

Abstract.
Assuming the four exponentials conjecture, Hansel and Safer showed that if a subset $S$ of the Gaussian integers is both $\alpha=-m+i$- and $\beta=-n+i$-recognizable, then it is syndetic, and they conjectured that $S$ must be eventually periodic. Without assuming the four exponentials conjecture, we show that if $\alpha$ and $\beta$ are multiplicatively independent Gaussian integers, and at least one of $\alpha$, $\beta$ is not an $n$-th root of an integer, then any $\alpha$- and $\beta$-automatic configuration is eventually periodic; in particular we prove Hansel and Safer’s conjecture. Otherwise, there exist non-eventually periodic configurations which are $\alpha$-automatic for any root of an integer $\alpha$. Our work generalises the Cobham-Semenov theorem to Gaussian numerations.
Key words and phrases:
automata sequences; Gaussian numeration; Cobham’s theorem
2020 Mathematics Subject Classification:
11B85, 13F07, 68Q45
The first author was supported by ANID/FONDECYT Postdoctorado 3230159 (year 2023).
The third author was supported by the EPSRC, grant number EP/V007459/2.
1. Introduction
A sequence $a=(a_{n})_{n\geq 0}$ is $k$-automatic if $a_{n}$ is a finite-state function of the base-$k$ expansion of $n$. Let $k$ and $\ell$ be multiplicatively independent positive integers. Cobham’s theorem [9] tells us that a sequence is both $k$- and $\ell$-automatic if and only if it is eventually periodic. This theorem is arguably the most significant result concerning automatic sequences. Its importance is reflected in the fact that it is connected to many deep generalisations and proofs in different areas of mathematics, such as the work of Semenov and Muchnik in logic [24, 20], the work of Durand in dynamical systems and substitution theory [11, 12], the work of Bès in numeration [4], the work of Adamczewski and Bell in algebra [3, 1], and the work of the latter and also Schäfke and Singer in analysis and ODEs [2, 23]. This list is far from exhaustive.
While Cobham’s theorem has been extended in many directions, one notable gap is its extension from the integers to Euclidean domains. In this paper, we take a first step and extend Cobham’s theorem to the Gaussian integers. To formulate the theorem in this context, however, one must first establish what a $\beta$-automatic configuration is for a Gaussian integer $\beta$. In particular, this requires a numeration system with base $\beta$. The numeration system for the Gaussian integers with base $\beta=-1+i$ and digits 0,1 is described by Knuth [16, Section 4.1], who attributes it to Penney[22]. The only Gaussian integers for which any $z\in\mathbb{Z}[\mathrm{i}]$ has an expansion base-$\alpha$ with natural numbers as digits, are Gaussian integers $\alpha=-m\pm\mathrm{i}$, with digit set $\{0,1,\dots,m^{2}\}$, as shown by Kátai and Szabó in [15].
In [14], Hansel and Safer considered these Gaussian integers $\alpha=-m+\mathrm{i}$ as a base for Gaussian numeration systems ($m\geq 1$), with digit set $\{0,1,\dots,m^{2}\}$. Any two Gaussian integers of this form must be multiplicatively independent. Assuming the four exponentials conjecture, Hansel and Safer showed that for such Gaussian integers, if the characteristic function of an infinite $S\subset\mathbb{Z}[i]$ is both $\alpha$- and $\beta$-recognizable, then it is syndetic. They conjectured that this set must be eventually periodic, where we recall that a configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is eventually periodic if there exist $\mathbb{Z}$-linearly independent $p,q\in\mathbb{Z}[\mathrm{i}]$ such that $a_{z}=a_{z+p}=a_{z+q}$ whenever z[$\Fz\in\mathbb{Z}[\mathrm{i}]\backslash$ F for some finite set $F$.
As a result, this extension of Cobham was left unsettled because it was thought to depend on deep analytic number theory. This paper shows that it does not, surprisingly, and opens up the way to Cobham for general number rings. We prove Cobham’s theorem for the largest possible family of Gaussian numerations, with any generating finite digit set and without using the four exponentials conjecture. We show
Theorem 1.
Let $\alpha$ and $\beta$ be two multiplicatively independent Gaussian integers with $|\alpha|,|\beta|>1$. If one of $\alpha$ or $\beta$ is not the root of an integer, then a configuration is $\alpha$- and $\beta$-automatic if and only if it is eventually periodic.
In particular, Hansel and Safer’s conjecture is proved true, since m+i-m+i is not the root of an integer if $m>1$. The requirement that one of $\alpha$ and $\beta$ must be a root of an integer cannot be relaxed, as the characteristic function of $\mathbb{N}$ can be shown to be $\alpha$-automatic whenever $\alpha$ is a root of an integer. Multiplicatively independent Gaussian roots of integers will however obey a Semenov-type theorem [24, 20], see also the exposition [6], behaving like higher dimensional automatic configurations generated by a base-$k$ numeration system for $\mathbb{N}\times\mathbb{N}$, where projections of $\alpha$- and $\beta$-automatic configurations are eventually periodic along any horizontal or vertical direction.
We base our proof on Thijmen Krebs’ alternative proof of Cobham’s theorem[17]. In Section 2, we provide the required background on Gaussian numerations, automaticity and Dirichlet approximation in $\mathbb{Z}[\mathrm{i}]$. In Section 3, we use automaticity and the pumping lemma to generate 2 linearly independent periods, provided that one of the bases is not a root of an integer. In Section 4, we discuss two closely related notions of periodicity that we will use, and a version of Fine and Wilf’s theorem that will guarantee that local periods can propagate to global periods. Finally in Section 5 we prove our main result.
2. Preliminaries
2.1. Gaussian numerations
We assume the reader is familiar with the fact that the Gaussian integers form a Euclidean domain. In what follows, we discuss some basic facts about the Gaussian integers and their numeration systems that shall be of use to us. For a more in-depth study of Gaussian numeration systems, we direct the reader to [18, Chapter 7].
For $z\in\mathbb{Z}[\mathrm{i}]$, we let $|z|$ denote the Euclidean norm of $z$. Given a Gaussian integer $\gamma\in\mathbb{Z}[\mathrm{i}]$ with $\lvert\gamma\rvert>1$ and a finite subset $D\subset\mathbb{Z}[\mathrm{i}]$ containing 0, to each word $w=w_{n-1}\dotsc w_{0}\in D^{*}$ we associate the number
$w]_{\gamma}=\sum_{j=0}^{n-1}w_{j}\gamma^{j};$
note that $w_{0}$ is the least significant digit. This notation extends to sets of words $L\subseteq D^{*}$, by writing $L]_{\gamma}=\{[w]_{\gamma}\>:\>w\in L\}$. We say that $w\in D^{*}$ is a $(\gamma,D)$-expansion of $z\in\mathbb{Z}[\mathrm{i}]$ if $w]_{\gamma}=z$. We say that $(\gamma,D)$ is an integral numeration system if every Gaussian integer $z$ has a unique expansion $n=\sum_{j=0}^{k_{n}-1}d_{j}\gamma^{j}$ with $d_{j}\in D$. In this case we write $(n)_{\gamma}=d_{k}\dots d_{0}$. A set $D\in\mathbb{Z}[\mathrm{i}]$ is a complete residue system for $\mathbb{Z}[\mathrm{i}]\pmod{\gamma}$ if every Gaussian $z$ is congruent modulo $\gamma$ to a unique element of $D$. We record the following result [18, Theorems 7.4.1, 7.4.2, 7.4.3], which indicates that we have an abundance of Gaussian numeration systems.
Theorem 2.
Let $|\gamma|>1$. Let $0\neq\gamma=a+bi\in\mathbb{Z}[\mathrm{i}]$, and let $\lambda=\gcd(a,b)$. Then:
  1. (1)
    The set
    $D(\gamma)=\biggl\{p+\mathrm{i}q:p=0,1,\dots,\frac{|\gamma|^{2}}{\lambda}-1,q=0,1,\dots,\lambda-1\biggr\}$
    is a complete residue system for $\mathbb{Z}[\mathrm{i}]\bmod\gamma$.
  2. (2)
    Let $D$ be a finite set in $\mathbb{Z}[\mathrm{i}]$. If every Gaussian integer has an expansion using $(\gamma,D)$, then $(\gamma,D)$ is a numeration system if and only if $D$ is a complete residue system for $\mathbb{Z}[\mathrm{i}]\mod\gamma$ and $0\in D$.
  3. (3)
    [10] Let $|\gamma|\geq\sqrt{5}$ and let
    $D=\biggl\{d\in\mathbb{Z}[\mathrm{i}]:\operatorname{Re}\Bigl(\frac{d}{\gamma}\Bigr),\operatorname{Im}\Bigl(\frac{d}{\gamma}{}\Bigr)\in\Bigl[-\frac{1}{2},\frac{1}{2}\Bigr)\biggr\}.$
    Then $(\gamma,D)$ is a numeration system.
  4. (4)
    [15] Let $N\coloneq|\gamma|^{2}$ and let
    $D=\{0,1,\dots,N-1\}.$
    Then $(\gamma,D)$ is a numeration system for $\mathbb{Z}[\mathrm{i}]$ if and only if $\gamma=-n\pm\mathrm{i}$.
2.2. Dirichlet approximation in $\mathbb{Z}[\mathrm{i}]$
In his original proof, Cobham needed the fact that for natural numbers $\alpha,\beta\geq 2$, the multiplicative group generated by the natural $\alpha$ and $\beta$ is dense in the positive real line. It turns out that this result is very hard to generalise to $\mathbb{C}$. Hansel and Safer showed that if $\alpha$ and $\beta$ are multiplicatively independent Gaussians, then the four exponentials conjecture implies that the multiplicative group $\langle\alpha,\beta\rangle^{\times}=\{\alpha^{m}\beta^{n}\>:\>m,n\in\mathbb{Z}\}$ is dense in $\mathbb{C}$. In our proof of Theorem 1, we only need that there are no isolated points in $\langle\alpha,\beta\rangle^{\times}$, a much weaker result.
Lemma 3 ([5], Lemma 3).
Let $\alpha,\beta\in\mathbb{Z}[\mathrm{i}]$ be two multiplicatively independent Gaussian integers with $\lvert\alpha\rvert,\lvert\beta\rvert>1$. Then $1$ is not an isolated point in the multiplicative group $\langle\alpha,\beta\rangle^{\times}$.
Proof.
We need to prove that $\alpha$ and $\beta$ are multiplicatively dependent if $1$ is isolated in $G=\langle\alpha,\beta\rangle^{\times}$. Suppose that $d\in\mathbb{C}\setminus\{0\}$ is a density point of $G$, say that it is the limit of $g_{n}\in G$. Then $g_{n+1}/g_{n}$ converges to $1$, and since $1$ is isolated this means that the sequence is eventually constant. It follows that $G$ is a discrete subset of the punctured plane if $1$ is isolated. Let $H\subset G$ be the cyclic subgroup that is generated by $\alpha$. Every residue class of the quotient $G/H$ has a representative in the annulus $A=\{z:1\leq|z|\leq|\alpha|\}$. Since $G\cap A$ is compact and discrete, it is finite. Therefore, $G/H$ is finite and $\beta^{n}\in H$ for some $n$, which means that $\alpha$ and $\beta$ are multiplicatively dependent. ∎
Corollary 4.
Let $\varepsilon>0$. Given two non-zero multiplicatively independent Gaussian integers $\alpha,\beta\in\mathbb{Z}[\mathrm{i}]$ with $\lvert\alpha\rvert,\lvert\beta\rvert>1$, there exist $m,n>0$ such that
(1) $\lvert\alpha^{m}-\beta^{n}\rvert<\varepsilon\lvert\beta^{n}\rvert$.
Proof.
Without loss of generality, we may assume that $\lvert\alpha\rvert>\lvert\beta\rvert$, as the only Gaussian integers of absolute value $1$ are the units, and thus two Gaussian integers of the same absolute value must be multiplicatively dependent. Since $1$ is not an isolated point in $\langle\alpha,\beta\rangle^{\times}$, there exist infinitely many values $m,n\in\mathbb{Z}$ for which $\lvert\alpha^{m}\beta^{-n}-1\rvert<\varepsilon$. Furthermore, $m$ and $n$ must have the same sign for all sufficiently small $\varepsilon$, since otherwise $\lvert\alpha^{m}\beta^{-n}-1\rvert$ is bounded from below by $\min(\lvert\alpha\beta\rvert-1,1-\lvert\alpha\beta\rvert^{-1})>0$.
Once again, without loss of generality, we may assume for sufficiently small $\varepsilon$ that $m>0$ due to the continuity of the function $z^{-1}$ near $1$. Thus, by multiplying our inequality by $\lvert\beta^{n}\rvert$, we obtain (1). ∎
Recall that $\alpha,\beta\in\mathbb{Z}[\mathrm{i}]$ are multiplicatively independent if there are no natural numbers $j,k$ such that $\beta^{j}=\alpha^{k}$. It takes some effort to decide whether numbers are multiplicatively dependent. The following classical result, appears as Corollary 3.12 in [21] and is known as Niven’s theorem.
Theorem 5.
Let $\alpha\in\mathbb{Z}[i]$ be such that $\alpha^{j}\in\mathbb{Z}$ for some $j\in\mathbb{N}$. Then $\text{arg}(z)$ is a multiple of $\pi/4$.
Hansel and Safer considered basis $\alpha=-m+i$ for $m\geq 1$. It turns out that the only numbers that are multiplicatively dependent on $\alpha$ are powers of $\alpha$.
Theorem 6.
Let $\alpha=-m+i$ for $m\geq 1$. If $\beta^{j}=\alpha^{k}$ for $j,k\in\mathbb{N}$, then, up to a unit, $\beta$ is a power of $\alpha$.
Proof.
Let $y=N(\beta)$, the norm of $\beta$. Taking norms we find $y^{j}=(m^{2}+1)^{k}$. Therefore, $j\cdot\text{ord}_{p}(y)=k\cdot\text{ord}_{p}(m^{2}+1)$ for all (rational) primes $p$. By dividing out common divisors, we may suppose that $\text{gcd}(j,k)=1$. It follows that $k$ divides all $\text{ord}_{p}(y)$. In particular, $y=z^{k}$ for some $z$ such that $z^{j}=m^{2}+1$. By Mihăilescu’s theorem [19], $j=1$. Since we divided out a common factor $h$, it follows that $\beta^{h}=\alpha^{hk}$. Up to a root of unity, $\beta$ is equal to $\alpha^{k}$. ∎
Even the seemingly simple task of deciding whether $\alpha$ and $\beta$ are multiplicatively dependent runs into a weak form of Catalan’s conjecture, wherein one of the powers is a square. Such number theoretic difficulties naturally arise when considering the extension of Cobham’s theorem from the integers to other domains. This may explain why this extension has been neglected so far.
2.3. Finite automata and automaticity
A deterministic finite automaton with output (DFAO) is defined by a $6$-tuple $\mathcal{A}=(S,D,\delta,s_{0},A,\tau)$, where $S$ is a finite set of states, $s_{0}\in S$ is the initial state, $D$ and $A$ are finite sets called the input and output alphabet, respectively, $\delta\colon S\times D\to S$ is the transition function and $\tau\colon S\to A$ is the output function. The transition function extends to the set $D^{*}$ of all words with letters in $S$ via the recurrence relation:
$\delta(s,wa)=\delta(\delta(s,w),a),\quad w\in D^{*},a\in D,$
and thus every DFAO determines a function $D^{*}\to A$ given by $w\mapsto\tau(\delta(s_{0},w))$. We shall use DFAOs by feeding in corresponding expansions of Gaussian integers in terms of powers of $\gamma$; when we want to emphasize this connection between the automaton and the associated $\gamma\in\mathbb{Z}[\mathrm{i}]$, we include appropriate subscripts in the corresponding sets, say, $\mathcal{A}_{\gamma}=(S_{\gamma},D_{\gamma},\delta_{\gamma},s_{0,\gamma},A_{\gamma},\tau_{\gamma})$.
We will need to consider automatic configurations defined by DFAOs with redundant digit sets, i.e. Gaussian numerations where each Gaussian integer has at least one expansion.
Definition 7.
Let $\gamma\in\mathbb{Z}[\mathrm{i}]$ be a Gaussian integer with $|\gamma|>1$ and let $D\subset\mathbb{Z}[\mathrm{i}]$ be a finite subset for which every $z\in\mathbb{Z}[\mathrm{i}]$ has at least one (finite) expansion with digits from $D$. Let $\mathcal{A}=(S,D,\delta,s_{0},A,\tau)$ be a DFAO. A configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is $(\gamma,D)$-automatic if $a_{z}=\tau(\delta(s_{0},w))$ whenever $w]_{\gamma}=z$.
Examples of $\beta$-automatic configurations include characteristic functions of $\beta$-recognizable sets, also called $\beta$-automatic sets. Any automatic configuration can be seen as a partition of $\mathbb{Z}[\mathrm{i}]$ into automatic sets.
Figure 1. A $(1+2\mathrm{i})$-automatic configuration with digit set $D=\{0,\pm 1,\pm\mathrm{i}\}$.
In this paper, almost all DFAOs are direct reading, that is, a word which represents $z$ is fed into the machine starting with its most significant digit. One exception is when we use Lemma 9; this will not be an issue as a configuration is $(\gamma,D)$-automatic in direct reading if and only if it is $(\gamma,D)$-automatic in reverse reading.
A set $S\subset\mathbb{Z}[\mathrm{i}]$ is $(\gamma,D)$-automatic if its indicator function is $(\gamma,D)$-automatic. In what follows, we shall assume that $\alpha,\beta\in\mathbb{Z}[\mathrm{i}]$ are two multiplicatively independent Gaussian integers, and we shall use the letter $\gamma$ to refer generically to either of them.
Remark 8.
We say that a DFAO $\mathcal{A}_{\gamma}$ is consistent if, whenever $w_{1}]_{\gamma}=[w_{2}]_{\gamma}$, then $\tau_{\gamma}(\delta_{\gamma}(s_{0,\gamma},w_{1}))=\tau_{\gamma}(\delta_{\gamma}(s_{0,\gamma},w_{2}))$. If $(\gamma,D)$ is integral, then any $(\gamma,D)$-automaton is consistent. Note that Definition 7 implicitly requires that the DFAO defining a $(\gamma,D)$-automatic configuration is consistent. Henceforth all DFAOs we work with are assumed consistent.
The classical definition of automaticity uses an integral numeration system $(\gamma,D)$. Proposition 10 below tells us that if $(\gamma,D)$ is integral and $D\subset D^{\prime}$, then the family of $(\gamma,D)$-automatic configurations equals the family of $(\gamma,D^{\prime})$-automatic configurations generated by consistent automata. For this reason, if $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is $(\gamma,D)$-automatic, then we will sometimes write that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is $\gamma$-automatic if we do not need to state what the digits are.
The following is based on [18, Lemma 7.1.1]; we include it for completeness.
Lemma 9.
Let $|\beta|>1$, let $(\beta,D)$ be an integral Gaussian numeration and let $D\subset D^{\prime}$. Then there exists a 1-uniform transducer, with a terminal function that sends any $(\beta,D^{\prime})$ expansion of $z$ to $(z)_{D,\beta}\in D^{*}$.
Proof.
The idea is as follows. We go through each digit of an expansion $d_{n}^{\prime}\dotsc d_{0}^{\prime}\in(D^{\prime})^{*}$ for a Gaussian integer $z\in\mathbb{Z}[\mathrm{i}]$, starting from the least significant digit $d_{0}^{\prime}$, which we replace by the unique digit $d_{0}\in D$ which is congruent to $d_{0}^{\prime}$ modulo $\beta$. We now have a carry $c_{0}$ for which the equality $\beta c_{0}=d_{0}^{\prime}-d_{0}$ holds, so $z=[d_{n}^{\prime}\cdots d_{1}^{\prime}d_{0}]_{\beta}+\beta c_{0}$. We iterate this process, at each step defining $c_{k}=d_{k}^{\prime}-d_{k}$ and $d_{k+1}$ as the only element of $D$ congruent to $c_{k}+d_{k+1}^{\prime}$, so that $d_{k+1}^{\prime}+c_{k}=\beta c_{k+1}+d_{k+1}$ for some value $c_{k+1}\in\mathbb{Z}[\mathrm{i}]$, which will become the next carry. At each step the equality $z=[d_{n}^{\prime}\dotsc d_{k+1}^{\prime}d_{k}\dotsc d_{0}]_{\beta}+\beta^{k}c_{k}$ holds. After we modify all of the digits $d_{0}^{\prime},\dotsc,d_{n}^{\prime}$, we have an expression of the form $z=\beta^{n}c_{n}+[d_{n}\dotsc d_{0}]_{\beta}$, so we only need to append the digits of the $(\beta,D)$-expansion of $c_{n}$ at the start to get a full $(\beta,D)$-expansion of $z$.
For this procedure to work properly, at each step we need to keep track of the value of the carry $c_{k}$. It can be seen that, if we start with an initial carry of 0, then there are finitely many choices for each $c_{k}$, and so, this process can be carried out by a transducer, whose set of states corresponds to all possible carries. Indeed, set $m=\max\{\lvert d^{\prime}-d\rvert:d^{\prime}\in D^{\prime},d\in D\}$ and $c=m/(|\beta|-1)$. Note that if $|c_{k}|<c$, then $|c_{k+1}|\leq(|c_{k}|+|d-d^{\prime}|)/|\beta|<\left|(c+m)/\beta\right|=c$. This shows that the state set $Q$ is finite.
Formally, consider the subsequential finite transducer $\mathcal{A}=(Q,D^{\prime},\delta,q_{0},D,\tau)$ defined as follows: $Q=\{s\in\mathbb{Z}[\mathrm{i}]:|s|<c\}$ is the set of possible carries, and the set of edges (which defines $\delta$ and $\tau$) is
$\{s\stackrel{{\scriptstyle d^{\prime}/d}}{{\longrightarrow}}s^{\prime}:s+d^{\prime}=\beta s^{\prime}+d\}.$
Let $d_{n}^{\prime}\dots d_{0}^{\prime}\in D^{\prime*}$ be a $(\beta,D^{\prime})$-expansion of $z$. Starting at $s_{0}=0$ there is a unique path
$s_{0}\stackrel{{\scriptstyle d^{\prime}_{0}/d_{0}}}{{\longrightarrow}}s_{1}\stackrel{{\scriptstyle d^{\prime}_{1}/d_{1}}}{{\longrightarrow}}s_{2}\stackrel{{\scriptstyle d^{\prime}_{2}/d_{2}}}{{\longrightarrow}}\dots\stackrel{{\scriptstyle d^{\prime}_{n-1}/d_{n-1}}}{{\longrightarrow}}s_{n}\stackrel{{\scriptstyle d^{\prime}_{n}/d_{n}}}{{\longrightarrow}}s_{n+1}.$
Define the terminal function $t:Q\rightarrow D^{*}$ as $t(s)=(s)_{(\beta,D)}$. It can be verified that $z=\sum_{j=0}^{n}d_{j}\beta^{j}+s_{n+1}\beta^{n+1}$, hence the output of the transducer on inputting $d_{n}^{\prime}\dots d_{0}^{\prime}\in D^{\prime*}$ is
$t(s_{n+1})d_{n}\dots d_{0}=(z)_{(\beta,D)}.$
The following proposition is reminiscent of Cabezas and Leroy’s [7, Proposition 4.1], which gives a similar result in the primitive, aperiodic, substitutional case.
Proposition 10.
Let $|\beta|>1$, let $(\beta,D)$ be an integral Gaussian numeration and let $D\subset D^{\prime}$. Then $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is $(\beta,D)$-automatic if and only if it is $(\beta,D^{\prime})$-automatic and generated by a consistent automaton.
Proof.
Let $(Q,D^{\prime},\delta,q_{0,T},D,\tau,t)$ be the transducer with terminal function $t$ as defined in Lemma 9, and let $(Q_{D},D,\delta_{D},q_{0,D},A_{D},\tau_{D})$ be the $(\beta,D)$-automaton that generates the $(\beta,D)$-automatic configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$. Construct a $(\beta,D^{\prime})$-automaton $\mathcal{A}=(Q\times Q_{D},D^{\prime},\delta^{\prime},[q_{0,D},q_{0,T}],A_{D},\tau^{\prime})$ by defining $\delta^{\prime}:Q\times Q_{D}\times D^{\prime}\rightarrow Q\times Q_{D}$ and $\tau^{\prime}:Q_{D}\rightarrow A_{D}$ as
$\delta^{\prime}([p,q],d)=[\delta(p,\tau(d)),\tau(q,d)]\mbox{ and }\tau^{\prime}([p,q])=\tau_{D}(p).$
One uses induction to verify that $\mathcal{A}$ generates $(a_{z})$ in $(\beta,D^{\prime})$-reading. Note furthermore that the automaton $\mathcal{A}$ constructed in Proposition 10 is consistent.
Conversely, if $\mathcal{A}$ is a consistent automaton generating the $(\beta,D^{\prime})$-automatic configuration $(a_{z})_{z\in\mathbb{Z}[i]}$, then one can restrict to $D$-inputs to see that the configuration is also $(\beta,D)$-automatic. ∎
Let $B(z,r)\subset\mathbb{C}$ denote the ball centred at $z$ with radius $r$. Next we choose an appropriate set $D_{\gamma}$ which by the next lemma will be large enough to ensure that every Gaussian integer relatively close to 0 has a short $(\gamma,D_{\gamma})$-expansion.
Lemma 11.
Let $(\gamma,D)$ be an integral numeration system. Let $r\geq 2$ be chosen so that
(2) $D\subset D_{\gamma}\coloneq B(0,r\lvert\gamma\rvert)\cap\mathbb{Z}[\mathrm{i}].$
If $z\in\mathbb{Z}[\mathrm{i}]$ satisfies $\lvert z\rvert\leq r\lvert\gamma\rvert^{n}$, then there exists a word $w=d_{n-1}\dotsc d_{0}\in D_{\gamma}^{n}$ for which $w]_{\gamma}=z$, i.e. $z$ has at least one $(\gamma,D_{\gamma})$-expansion of length at most $n$.
Proof.
We proceed by induction on $n$; the result is evidently true if $\lvert z\rvert\leq r\lvert\gamma\rvert$ by our choice of $D_{\gamma}$. Suppose that the result is valid for all Gaussian integers in $B(0,r\lvert\gamma\rvert^{n})\cap\mathbb{Z}[\mathrm{i}]$, and let $z\in B(0,r\lvert\gamma\rvert^{n+1})\cap\mathbb{Z}[\mathrm{i}]$. We have $\lvert\frac{z}{\gamma}\rvert\leq\frac{r\lvert\gamma\rvert^{n+1}}{\lvert\gamma\rvert}=r\lvert\gamma\rvert^{n}$; we claim that this implies that there exists a Gaussian integer $q\in\mathbb{Z}[\mathrm{i}]$ with $|q|\leq r\lvert\gamma\rvert^{n}$ and $|q-\frac{z}{\gamma}|\leq\sqrt{2}$. Suppose first that $\frac{z}{\gamma}=u_{1}+u_{2}\mathrm{i}$ with $u_{1},u_{2}\in\mathbb{R}_{\geq 0}$ and set $q=\lfloor u_{1}\rfloor+\lfloor u_{2}\rfloor\mathrm{i}$. Then, $\lvert q\rvert^{2}=\lfloor u_{1}\rfloor^{2}+\lfloor u_{2}\rfloor^{2}\leq u_{1}^{2}+u_{2}^{2}\leq(r\lvert\gamma\rvert^{n})^{2}$. The proof for the other three quadrants proceeds in the same fashion, replacing $\lfloor u_{j}\rfloor$ by $\lfloor-u_{j}\rfloor$ when that coordinate is negative. The inequality $\lvert q-\frac{z}{\gamma}\rvert\leq\sqrt{2}$ follows immediately from the fact that $\lvert x-\lfloor x\rfloor\rvert<1$.
[…]
$z=q\gamma+(z-q\gamma)=d_{n-1}\gamma^{n}+\cdots+d_{0}\gamma+d^{*},$
where all digits $d_{0},\dotsc,d_{n-1},d^{*}\in D_{\gamma}$, hence the word $d_{n-1}\dotsc d_{0}d^{*}\in D_{\gamma}^{n+1}$ is a $(\gamma,D_{\gamma})$-expansion of $z$ of length $n+1$. ∎
3. Generating periods
In this section, in Lemma 12 we exploit the structure of $(\gamma,D)$-expansions of Gaussian integers to show repetitive behaviour on a corresponding automatic configuration. Lemmas 14 and 15 will allow us to later create pairs of non-collinear vectors which we will bootstrap, to show that configurations satisfying the conditions of Theorem 1 are eventually periodic.
Let $\gamma$ be a Gaussian integer with $\lvert\gamma\rvert>1$, and $D$ a digit set for which each $z\in\mathbb{Z}[\mathrm{i}]$ has at least one $(\gamma,D)$-expansion. Suppose that $\mathcal{A}_{\gamma}=(S,D,\delta,s_{0},A,\tau)$ is a consistent $(\gamma,D)$-automaton, and let $s\in S$; we define:
(3) $L_{\gamma,s}\coloneqq\{w\in D^{*}\>:\>\delta(s_{0},w)=s\}.$
Thus, the set $L_{\gamma,s}]_{\gamma}$ corresponds to all the Gaussian integers which have at least one $(\gamma,D)$-expansion that ends at state $s$ in the automaton $\mathcal{A}$. The sets $\{[L_{\gamma,s}]_{\gamma}\>:\>s\in S\}$ are a cover of $\mathbb{Z}[\mathrm{i}]$, but they might not be a partition as some Gaussian integers might have more than one $(\gamma,D)$-expansion.
Lemma 12.
Let $\mathcal{A}=(S,D,\delta,s_{0},A,\tau)$ be a $(\gamma,D)$-consistent DFAO, and let $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ be the corresponding automatic configuration. Then, for every $x,y\in[L_{\gamma,s}]_{\gamma}$, any $n\in\mathbb{N}$ and every $z\in[D^{n}]_{\gamma}$, we have the following equality:
(4) $a_{x\gamma^{n}+z}=a_{y\gamma^{n}+z},$
that is, this identity holds for any $x$ and $y$ with $(\gamma,D)$-expansions ending at the same state of $\mathcal{A}$, and any $z$ that has a $(\gamma,D)$-expansion of length $n$ or less.
Proof.
By definition, there exist $w_{1},w_{2}\in L_{\gamma,s}$ such that $w_{1}]_{\gamma}=x,[w_{2}]_{\gamma}=y$. Thus, for any $v\in D^{n}$, we have $\delta(s_{0},w_{j}v)=\delta(\delta(s_{0},w_{j}),v)=\delta(s,v)$. Hence, as $w_{j}v]_{\gamma}=[w_{j}]_{\gamma}\gamma^{n}+[v]_{\gamma},j\in\{1,2\}$, if we write $z=[v]_{\gamma}$, we have
$a_{x\gamma^{n}+z}=\tau(\delta(s_{0},w_{1}v))=\tau(\delta(s,v))=\tau(\delta(s_{0},w_{2}v))=a_{y\gamma^{n}+z},$
since $w_{1}v$ and $w_{2}v$ are $(\gamma,D)$-expansions for $x\gamma^{n}+z$ and $y\gamma^{n}+z$, respectively. ∎
[…]
Let $S_{\infty}$ be the set of states in $S_{\beta}$ for which $L_{\beta,s}]_{\beta}$ is infinite, where $L_{\beta,s}$ is defined as in (3). We will need $S_{\infty}$ to contain certain triples of non-collinear points. The following lemma tells us that this is guaranteed if $\beta$ is not an $n$-th root of an integer. We use the existence of a cycle rooted at $s\in S_{\infty}$ to deduce that Gaussian integers with expansions ending at $s$ do not live in a one dimensional subspace unless $\beta$ is restricted. The existence of such a cycle is in general not guaranteed, but the pumping lemma can be used to create a weaker version of this condition which will be sufficient. For a word $w$ over an alphabet $A$ we let $|w|$ denote the length of $w$. (It will be clear from the context that the object considered is a word $w$ and not a Gaussian integer $z$.)
Definition 13.
Let $\mathcal{A}_{\beta}$ be a DFAO, and let $s\in S_{\infty}$. Then, by the pigeonhole principle, there exist $s^{\prime}\in S_{\infty}$ and three words $w_{1},w_{2},w_{3}\in D_{\beta}^{*}$ for which $\delta_{\beta}(s_{0},w_{1})=s^{\prime}$, $\delta_{\beta}(s^{\prime},w_{2})=s^{\prime}$ and $\delta_{\beta}(s^{\prime},w_{3})=s$. (If there is a cycle in the underlying graph of $\mathcal{A}_{\beta}$ rooted at $s$, we can take $s=s^{\prime}$ and $w_{3}$ to be the empty word.) Furthermore, since $s^{\prime}\in S_{\infty}$, there are infinitely many choices for $w_{1}]_{\beta}$, so we pick one so that $w_{1}]_{\beta}\cdot(1-\beta^{\lvert w_{2}\rvert})\neq[w_{2}]_{\beta}$. Note that for each $j$, $u_{j}=[w_{1}(w_{2})^{j}w_{3}]_{\beta}$ is a word whose $(\beta,D_{\beta})$-expansion ends at state $s$. We call such numbers $u_{j}$ return numbers to $s$.
Lemma 14.
Let $s\in S_{\infty}$, and suppose that $(u_{j})$ are return numbers to $s$. If there exists an arithmetic progression of indices such that $u_{k},u_{k+\ell},u_{k+2\ell}$ are collinear return numbers, then $\beta$ is the root of an integer.
Proof.
We have $u_{j}=[w_{1}w_{2}^{j}w_{3}]_{\beta}$ where the words $w_{i}$ are as in Definition 13. First suppose that $s=s^{\prime}$ and $w_{3}$ is the empty word, i.e., that there is a cycle rooted at $s$.
Let $N=\lvert w_{2}\rvert$ and $h=[w_{2}]_{\beta}$. The sequence of return numbers $(u_{j})_{j\in\mathbb{N}}$ satisfies the recurrence relation $u_{j+1}=\beta^{N}u_{j}+h$, from which the following equalities immediately follow:
$u_{k+\ell}$ $=\beta^{N\ell}u_{k}+h(\beta^{N(\ell-1)}+\beta^{N(\ell-2)}+\cdots+\beta^{N}+1)$
$u_{k+2\ell}$ $=\beta^{N\ell}u_{k+\ell}+h(\beta^{N(\ell-1)}+\beta^{N(\ell-2)}+\cdots+\beta^{N}+1)$
${}\implies u_{k+2\ell}-u_{k+\ell}$ $=\beta^{N\ell}(u_{k+\ell}-u_{k}).$
We now face two scenarios. If $u_{k+\ell}-u_{k}\neq 0$, we may divide by this term and obtain the equality $\beta^{N\ell}=\frac{u_{k+2\ell}-u_{k+\ell}}{u_{k+\ell}-u_{k}}$. If $u_{k},u_{k+\ell},u_{k+2\ell}$ lie on the same line in the plane, then $u_{k+2\ell}-u_{k+\ell}$ and $u_{k+\ell}-u_{k}$ must lie on a parallel line which passes through the origin, i.e. there exists a real constant $\lambda$ for which $u_{k+2\ell}-u_{k+\ell}=\lambda(u_{k+\ell}-u_{k})$. But then $\beta^{N\ell}=\lambda\in\mathbb{R}$, which is only possible if $\arg(\beta)$ is a rational multiple of $2\pi\mathrm{i}$.
Otherwise, if $u_{k+\ell}-u_{k}=0$, we obtain that $u_{j+\ell}=u_{j}$ for all $j\geq k$ from the aforementioned recurrence relation, so the sequence $(u_{j})_{j\in\mathbb{N}}$ remains bounded. Write $\hat{u}_{j}=u_{j}-\frac{h}{1-\beta^{N}}$; from this definition, it holds that:
$\hat{u}_{j+1}$ $=u_{j+1}-\frac{h}{1-\beta^{N}}$
$=\beta^{N}u_{j}+h-\frac{h}{1-\beta^{N}}$
$=\beta^{N}\left(\hat{u}_{j}+\frac{h}{1-\beta^{N}}\right)+h\left(1-\frac{1}{1-\beta^{N}}\right)$
$=\beta^{N}\hat{u}_{j}+\frac{\beta^{N}h}{1-\beta^{N}}+\frac{h((1-\beta^{N})-1)}{1-\beta^{N}}$
$=\beta^{N}\hat{u}_{j},$
and since $\lvert\beta\rvert>1$, we have that $\lvert\hat{u}_{j}\rvert\to\infty$ unless $\hat{u}_{0}=u_{0}-\frac{h}{1-\beta^{N}}=0$. Since $u_{0}=[w_{1}]_{\beta}$ and $h=[w_{2}]_{\beta}$, this is impossible as we chose $w_{1}$ and $w_{2}$ so that $w_{1}]_{\beta}\cdot(1-\beta^{\lvert w_{2}\rvert})\neq[w_{2}]_{\beta}$. As the sequence $(u_{j})_{j\in\mathbb{N}}$ is bounded if, and only if, $(\hat{u}_{j})_{j\in\mathbb{N}}$ remains bounded as well, we get a contradiction, so this scenario never happens.
Consider now the general case when $s\neq s^{\prime}$. We define $M=\lvert w_{3}\rvert$, $h^{\prime}=[w_{3}]_{\beta}$, and for each $j$, $u_{j}=[w_{1}(w_{2})^{j}]_{\beta}$, which are return numbers to $s^{\prime}$. Then we set $v_{k}=\beta^{M}u_{k}+h^{\prime}$, which are return numbers to $s$, and each term $v_{k}$ is obtained from the term $u_{k}$ via a dilation followed by a translation. As these transformations preserve collinearity, the return numbers $v_{k},v_{k+\ell},v_{k+2\ell}$ lie on the same line if, and only if, the respective $u_{k},u_{k+\ell},u_{k+2\ell}$ are collinear as well, and thus an application of the first case above to the state $s^{\prime}$ shows us that $\arg(\beta)$ is a rational multiple of $2\pi$. ∎
Lemma 15.
Given an $\alpha$- and $\beta$-automatic configuration, let $s\in S_{\infty}\subset S_{\beta}$. If $\beta$ is not a root of an integer, then there exists a state $t=t(s)\in S_{\alpha}$ for which we may choose three non-collinear Gaussian integers $x_{s,t},y_{s,t},z_{s,t}\in[L_{\beta,s}]_{\beta}\cap[L_{\alpha,t}]_{\alpha}$. That is, each of the $x_{s,t},y_{s,t},z_{s,t}$ have both a $(\beta,D_{\beta})$-expansion ending at state $s$ in the $(\beta,D_{\beta})$ automaton, and an $(\alpha,D_{\alpha})$-expansion ending at state $t$ in the $(\alpha,D_{\alpha})$ automaton.
Proof.
We pick a sequence of return numbers $(u_{j})_{j\in\mathbb{N}}$ to $s$ satisfying the conditions of Definition 13. Define a partition $\{U_{t}\}_{t\in S_{\alpha}}$ of $\mathbb{N}$ satisfying the condition that, whenever $j\in U_{t}$, the Gaussian integer $u_{j}$ has an $(\alpha,D_{\alpha})$-expansion ending at state $t$ in $\mathcal{A}_{\alpha}$; such a partition always exists, as every Gaussian integer has at least one $(\alpha,D_{\alpha})$-expansion. By van der Waerden’s theorem [25], there exists at least one state $t(s)\in S_{\alpha}$ for which the set $U_{t(s)}$ contains an arithmetic progression $\{k,k+\ell,k+2\ell\}$ of length $3$. By Lemma 14, the trio $u_{k},u_{k+\ell},u_{k+2\ell}$ cannot lie on the same line, as otherwise $\arg(\beta)$ would be rational. Furthermore, as all three points lie in $U_{t(s)}$, they have at least one $(\alpha,D_{\alpha})$-expansion ending at state $t(s)$. If we define $x_{s,t(s)}=u_{k},y_{s,t(s)}=u_{k+\ell},z_{s,t(s)}=u_{k+2\ell}$, we are done.
4. Periodic patterns on $\mathbb{Z}[\mathrm{i}]$
We will use the Gaussian integers $x_{s,t},y_{s,t},z_{s,t}$ from the statement of Lemma 15 to define vectors which define two-linearly independent periods for an $\alpha$- and $\beta$-automatic configuration. Given two $\mathbb{Z}$-linearly independent Gaussian integers $p_{1}$, $p_{2}$, let $\langle p_{1},p_{2}\rangle=\{m_{1}p_{1}+m_{2}p_{2}:m_{i}\in\mathbb{Z}\}$ denote the lattice generated by them. We say that $z_{1},z_{2}\in\mathbb{Z}[\mathrm{i}]$ are congruent modulo $\langle p_{1},p_{2}\rangle$, written $z_{1}\equiv z_{2}\pmod{\langle p_{1},p_{2}\rangle}$, if $z_{1}-z_{2}\in\langle p_{1},p_{2}\rangle$.
Definition 16.
Let $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ be a configuration, let $U\subseteq\mathbb{Z}[\mathrm{i}]$, and let $p_{1},p_{2}\in\mathbb{Z}[\mathrm{i}]$ be $\mathbb{Z}$-linearly independent. We say that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has period $(p_{1},p_{2})$ in $U$ if for any $z_{1},z_{2}\in U$, whenever $z_{1}\equiv z_{2}\pmod{\langle p_{1},p_{2}\rangle}$, we also have $a_{z_{1}}=a_{z_{2}}$.
We say that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has step-period $(p_{1},p_{2})$ in $U$ if whenever $z$ and $z+p_{i}$ belong to $U$ for some $i$, then $a_{z}=a_{z+p_{i}}$.
It is straightforward to see that if a configuration has period $(p_{1},p_{2})$ in $U$, then it has step-period $(p_{1},p_{2})$ in $U$. The converse is true if $U$ is reasonable and we shrink it slightly, as shown in Lemma 19. In particular, for $U=\mathbb{Z}[\mathrm{i}]$, the two notions are equivalent. For an example of a set which is step-periodic but not periodic, see Figure 2.
Figure 2. A configuration with step-period $(1+4\mathrm{i},1-4\mathrm{i})$ on the ball of radius $5+\epsilon$, for some $\epsilon<1$, around the origin that does not satisfy the definition of period. The values z=5z=-5 and $z=1$ are congruent modulo the lattice $\langle 1+4\mathrm{i},1-4\mathrm{i}\rangle$, so the definition of periodicity requires them to be the same color; however, all four neighbours $5\pm(1\pm 4\mathrm{i})$ fall outside the ball, so step-periodicity is not enough to guarantee that they do actually have the same color.
Note that our notion of periodicity is different to that of Hansel and Safer in [14]. They define $S\subset\mathbb{Z}[\mathrm{i}]$ to be periodic if there exists $p\in\mathbb{Z}[\mathrm{i}]$ such that $s\in S$ if and only if $s+zp\in S$ for any $z\in\mathbb{Z}[\mathrm{i}]$, i.e., that $S$ is a union of cosets of an ideal $(p)$. We may even restrict to $p\in\mathbb{Z}$, since if $S$ is a union of cosets mod $(p)$, then it is also a union of cosets mod $(pq)$, and we may take $q$ to be the complex conjugate of $p$. Beyond the cosmetic difference that we define periodicity for configurations while they define periodicity for sets, if $U=\mathbb{Z}[\mathrm{i}]$, one can move between our definition and theirs. To see this, note that any lattice $\langle p_{1},p_{2}\rangle$, where $p_{1}$ and $p_{2}$ are $\mathbb{Z}$-linearly independent, contains elements $m$ and $n\mathrm{i}$ with $m$ and $n$ non-zero integers. Setting $p=\operatorname{lcm}(m,n)$, the lattice $\langle p_{1},p_{2}\rangle$ contains any element of the form $ap+bp\mathrm{i}$, i.e., it contains the ideal $(p)$. Conversely, being $p$-periodic in the sense of Hansel and Safer is just having periods $p$ and $p\mathrm{i}$ in our first definition.
The proof of the following lemma follows from the triangle inequality.
Lemma 17.
[…]
In what follows, we shall study how periodic behaviour extends from one set to another when they have a sufficiently large overlap. As our proofs deal with bounded sets, we have to take into account the fact that the behaviour might not be what is expected near the “border” of one such set, and thus we introduce the notation
$V^{\circ r}\coloneqq\{z\in V\>:\>B(z,r)\cap\mathbb{Z}[\mathrm{i}]\subseteq V\}.$
The following two lemmas are reminiscent of Fine and Wilf’s theorem [13], which in turn improves Bezout’s identity.
Lemma 18 (Period transfer lemma).
Let $U,V\subseteq\mathbb{Z}[\mathrm{i}]$ be two sets such that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has a step-period $(p,q)$ in $U$, and a period $(p^{\prime},q^{\prime})$ in $V$. Define the following values:
$\rho_{1}=\frac{1}{2}(\lvert p\rvert+\lvert q\rvert),\quad\rho_{2}=\max(\lvert p^{\prime}\rvert,\lvert q^{\prime}\rvert),\quad R=\rho_{1}+\rho_{2}.$
Suppose $U\cap V$ contains a ball of radius $R$. Then $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has a step-period $(p,q)$ in the set $U\cup V^{\circ\rho_{1}}$.
Proof.
Suppose both $z$ and $z+p$ belong to the set $U\cup V^{\circ\rho_{1}}$. We are going to show that $a_{z}=a_{z+p}$; this is immediate if both $z$ and $z+p$ are in $U$, so we can assume without loss of generality that either $z$ or $z+p$ belong to $V^{\circ\rho_{1}}$. Then, both of them belong to $V$; indeed, if we suppose $z\in V^{\circ\rho_{1}}$, we must have $z+p\in B(z,\rho_{1})\cap\mathbb{Z}[\mathrm{i}]\subseteq V$. The same argument holds for the second case.
Let $B^{*}$ be a ball of radius $R$ contained in $U\cap V$, and $B^{**}$ a ball of radius $\rho_{2}<R$ with the same centre. Note that $B^{**}$ is large enough to contain a complete set of representatives modulo $\langle p^{\prime},q^{\prime}\rangle$. Thus, as $z$ is in $V$, there exist $m,n\in\mathbb{Z}$ such that $z+mp^{\prime}+nq^{\prime}\in B^{**}\subseteq B^{*}$. Furthermore, the ball of radius $\rho_{2}$ centered at $z+mp^{\prime}+nq^{\prime}$ is contained in $B^{*}$, so $z+p+mp^{\prime}+nq^{\prime}\in B^{*}$ as well. Since $B^{*}\subseteq V$, $a_{z}=a_{z+mp^{\prime}+nq^{\prime}}$ and $a_{z+p}=a_{z+p+mp^{\prime}+nq^{\prime}}$ by $(p^{\prime},q^{\prime})$-periodicity at $V$. But $B^{*}$ is also contained in $U$, so $a_{z+mp^{\prime}+nq^{\prime}}=a_{(z+mp^{\prime}+nq^{\prime})+p}$. We conclude by transitivity that $a_{z}=a_{z+p}$. The argument for $q$ is similar. ∎
Lemma 19.
Let $U\subseteq\mathbb{Z}[\mathrm{i}]$ be a ball of radius $R$ such that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has step-period $(p,q)$ in $U$. Then $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has period $(p,q)$ in $U^{\circ r}$ for $r=|p|+|q|$.
Proof.
We start by defining the fundamental domain:
(5) $\Delta=\{Ap+Bq\>:\>0\leq A,B<1\}\cap\mathbb{Z}[\mathrm{i}].$
It is not hard to see that $\Delta$ is a complete set of representatives for the lattice $\langle p,q\rangle$, meaning that any element of $\mathbb{Z}[\mathrm{i}]$ is congruent to some element of $\Delta$.
Let $F\subset\mathbb{Z}[\mathrm{i}]$ be any connected subset of the Cayley graph of $(\mathbb{Z}[\mathrm{i}],+)$ with generating set $\{1,\mathrm{i}\}$, meaning that for any $u,v\in F$ there exists a sequence $u=u_{0},\dotsc,u_{n}=v$ of elements of $F$ satisfying $u_{j+1}-u_{j}\in\{\pm 1,\pm\mathrm{i}\}$. If we write $F^{(p,q)}\coloneqq\{ap+bq\>:\>a,b\in\mathbb{Z},a+b\mathrm{i}\in F\}$, then, for any elements $z,z^{\prime}$ in the Minkowski sum $F^{(p,q)}+\Delta$, we have that $z\equiv z^{\prime}\pmod{\langle p,q\rangle}$ if, and only if, there exists some sequence $z=z_{0},z_{1},\dotsc,z_{m}=z^{\prime}$ with $z_{j+1}-z_{j}\in\{\pm p,\pm q\}$ and every $z_{j}\in F^{(p,q)}+\Delta$, something that follows from the connectedness of $F$. Thus, if $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has step-period $(p,q)$ on $F^{(p,q)}+R$, whenever $z\equiv z^{\prime}\pmod{\langle p,q\rangle}$ we have $a_{z}=a_{z_{1}}=a_{z_{2}}=\cdots=a_{z_{m}}=a_{z^{\prime}}$, and hence $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has period $(p,q)$ in $F^{(p,q)}+\Delta$ as well, i.e. both properties are equivalent on any set of this form. Naturally, the same holds for any translate of a set of the form $F^{(p,q)}+\Delta$. We have shown that on a set of the form $F^{(p,q)}+\Delta$, step-periodicity is the same as periodicity.
Now, suppose that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has step-period $(p,q)$ in a ball $U$ of radius $R$. Then, it also has step-period $(p,q)$ in any subset of the form $F^{(p,q)}+\Delta$ contained in $U$, as defined above. Let $F$ be the largest connected subset of $\mathbb{Z}[\mathrm{i}]$ for which $F^{(p,q)}+\Delta\subseteq U$; as $r$ is at least the length of either diagonal of $\Delta$, then for any $z\in U^{\circ r}$ there exists some translate $z^{\prime}+\Delta$ that contains $z$ with $z^{\prime}\in F^{(p,q)}$. Thus, $U^{\circ r}\subseteq F^{(p,q)}+\Delta\subseteq U$. As the latter set has step-period $(p,q)$, the middle set has period $(p,q)$, as discussed above. This in turn implies that $U^{\circ r}$ has period $(p,q)$, being a subset of a set with this property. ∎
[…]
Lemma 20.
If $K\geq 28$, then the set $\{B(x\beta^{n},(1-\frac{8}{K})\lvert\beta^{n}\rvert)\>:\>x\in\mathbb{Z}[\mathrm{i}]\}$ is a cover of $\mathbb{C}$ (and thus of $\mathbb{Z}[\mathrm{i}]$).
Proof.
Geometrically, multiplication by $\beta^{n}$ corresponds to a rotation followed by a dilation by $\lvert\beta^{n}\rvert$. Thus, the result is equivalent to the set $\{B(x,1-\frac{8}{K})\>:\>x\in\mathbb{Z}[\mathrm{i}]\}$ being a cover of $\mathbb{C}$. Since the covering radius of the lattice $\mathbb{Z}[\mathrm{i}]$ is $\frac{\sqrt{2}}{2}$, this will be true for any value of $K$ that ensures that $1-\frac{8}{K}>\frac{1}{\sqrt{2}}$. Hence:
$\frac{8}{K}<\frac{2-\sqrt{2}}{2}\implies K>\frac{16}{2-\sqrt{2}}=8(2+\sqrt{2}).$
In particular, $28>8(2+\sqrt{2})$, so these balls cover $\mathbb{C}$ for any value of $K\geq 28$. ∎
5. Proof of the main theorem for $\mathbb{Z}[\mathrm{i}]$
Lemma 21.
Any eventually periodic configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is automatic for any numeration system $(\beta,D)$.
Proof.
We start by recalling some basic facts to ease the exposition. Given any positive integer $p$, every Gaussian integer $z\in\mathbb{Z}[\mathrm{i}]$ is congruent modulo $p$ to a unique element in the set $\{a+b\mathrm{i}\>:\>0\leq a,b<p\}$, which we shall denote by $z\bmod p$; this element may be computed by taking the remainder modulo $p$ of both the real and imaginary parts of $z$ (i.e. $a+b\mathrm{i}\bmod p=(a\bmod p)+(b\bmod p)\mathrm{i}$). Recall that for configurations supported on the whole of $\mathbb{Z}[\mathrm{i}]$ our notion of being periodic in Definition 16 is equivalent to the notion of $p$-periodicity introduced by Hansel and Safer, and furthermore we may assume $p$ to be real (more precisely, a positive integer); that is, a configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is periodic if, and only if, there exists $p\in\mathbb{Z}_{>0}$ for which $a_{z}=a_{z+kp}$ for any $z,k\in\mathbb{Z}[\mathrm{i}]$. We shall use those two facts to show that a fully periodic configuration is always $(\beta,D)$-automatic for all choices of $D$ and $\beta$; in what follows, we first assume that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is a periodic configuration, and $p\in\mathbb{Z}_{>0}$ is chosen such that this configuration is $p$-periodic in the sense of Hansel and Safer.
Note that the sequence $(\beta^{j}\bmod p)_{j\in\mathbb{N}}$ is eventually periodic; say that there exist $m>0$ and $n\geq 0$ such that $\beta^{j}\equiv\beta^{j+km}\pmod{p}$ for any $k\geq 0,j\geq n$. In particular, any $\beta^{j}$ for $j\geq n$ is congruent to some element of $\{\beta^{n},\beta^{n+1},\dotsc,\beta^{n+m-1}\}$. We take $n$ and $m$ to be the smallest possible values for these constants.
Define a deterministic finite automaton with state set $S=\{s_{z,j}\>:\>z=a+b\mathrm{i},0\leq a,b<p,0\leq j<n+m\}$, initial state $s_{0,0}$ and transition function given by:
$\delta(s_{z,j},\ell)=\begin{cases}s_{z+\beta^{j}\ell\bmod p,j+1}&\text{if }j<n+m-1,\\ s_{z+\beta^{j}\ell\bmod p,n}&\text{if }j=n+m-1.\end{cases}$
The eventual periodicity of $(\beta^{j}\bmod{p})$ shows that if $\delta(s_{0,0},w)=s_{z,j}$, then $\beta^{\lvert w\rvert}\equiv\beta^{j}\pmod{p}$. We claim that from this it follows that $w]_{\beta}\equiv z\pmod{p}$, if $w$ is input in reverse reading. The proof is by (structural) induction; the result is immediate when $\lvert w\rvert=1$, as $\delta(s_{0,0},\ell)=s_{\ell\bmod p,1}$. Thus, suppose that the result holds for a given $w\in D^{*}$, and let $\delta(s_{0,0},\ell w)=\delta(s_{z,j},\ell)=s_{z^{\prime},j^{\prime}}$; by the induction hypothesis, $w]_{\beta}\equiv z\pmod{p}$. Since $\beta^{\lvert w\rvert}\equiv\beta^{j}\pmod{p}$, we have:
$\ell w]_{\beta}=\beta^{\lvert w\rvert}\ell+[w]_{\beta}\equiv\beta^{j}\ell+z\equiv z^{\prime}\pmod{p}$,
which proves our claim. This result may be restated as the equality $\delta(s_{0,0},w)=s_{[w]_{\beta}\bmod{p},j}$, for some $j$ depending only on $\lvert w\rvert$.
As $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is $p$-periodic, we have that $z\equiv z^{\prime}\pmod{p}$ implies that $a_{z}=a_{z^{\prime}}$, so in particular $a_{z}=a_{z\bmod p}$ for any $z\in\mathbb{Z}[\mathrm{i}]$. If we define the output function as $\tau(s_{z,j})=a_{z}$, the corresponding reverse reading DFAO satisfies:
$\tau(\delta(s_{0,0},w))=\tau(s_{[w]_{\beta}\bmod p,j})=a_{[w]_{\beta}\bmod p}=a_{[w]_{\beta}},$
so this is a $(\beta,D)$-automaton that computes $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$.
If $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is eventually periodic, there exists a finite set $F\subset\mathbb{Z}[\mathrm{i}]$ and a fully periodic configuration $(b_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ such that $a_{z}=b_{z}$ for all $z\in\mathbb{Z}[\mathrm{i}]\setminus F$. Let $\tilde{F}\subset D^{*}$ be the set of all $w$ for which $w]_{\beta}\in D$. If $\tilde{F}$ is finite, then by adding finitely many states we can recognize all words of $\tilde{F}$, and define an output function given by $a_{[w]_{\beta}}$ whenever $w\in\tilde{F}$, and the original output $b_{[w]_{\beta}}$ otherwise. If $\tilde{F}$ is infinite (which may happen if $D$ is redundant), then we can use Proposition 10 to convert the numeration system to one where $\tilde{F}$ is finite. The result follows.
We can now prove our main theorem. The argument is an extension of Krebs’s argument [17] to the Gaussian numbers.
Theorem 1.
Let $\alpha$ and $\beta$ be two multiplicatively independent Gaussian integers with $|\alpha|,|\beta|>1$. Suppose that $\beta$ is not the root of an integer. Then a configuration is $\alpha$- and $\beta$-automatic if and only if it is eventually periodic.
Remark 22.
The statement of Theorem 1 is the best possible. For, suppose that $\alpha$ and $\beta$ are both roots of an integer. Then there exist configurations which are both $\alpha$- and $\beta$- automatic, but not eventually periodic. To see this we apply [5, Theorem 4], which tells us that $\mathbb{N}$ is a $\beta$-automatic set if and only if $\beta$ is a root of an integer. The characteristic function of $\mathbb{N}$ is not eventually periodic. It can be shown that a Semenov-Cobham type theorem exists for such bases.
Proof of Theorem 1.
By Lemma 21, any eventually periodic configuration $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is automatic for any base. Conversely, we shall use Proposition 10, which tells us that an automatic configuration generated by an integral numeration $(\gamma,D)$ is also automatic when generated by a consistent $(\gamma,D_{\gamma})$-automaton for a sufficiently enlarged digit set $D_{\gamma}$ satsifying the conditions of (2). Suppose that $f=(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ is $\alpha$- and $\beta$-automatic, generated by the finite state machines $\mathcal{A_{\alpha}}=(S_{\alpha},D_{\alpha},\delta_{\alpha},s_{0,\alpha},A_{\alpha},\tau_{\alpha})$ and $\mathcal{A_{\beta}}=(S_{\beta},D_{\beta},\delta_{\beta},s_{0,\beta},A_{\beta},\tau_{\beta})$ respectively, where $D_{\alpha}$ and $D_{\beta}$ are defined as in Equation (2). Let $\gamma\in\{\alpha,\beta\}$. Then by Lemma 11, if $z\in B(\gamma^{n},|\gamma^{n}|)\subset B(0,r|\gamma^{n}|)$, where $r\geq 2$, then it will have a short $(\gamma,D_{\gamma})$-expansion, specifically $B(\gamma^{n},|\gamma^{n}|)\subset[D_{\gamma}^{n}]_{\gamma}$. Taking $L_{\gamma s}$ as defined in (3), by Lemma 12
$a_{x\gamma^{n}+z}=a_{y\gamma^{n}+z}$
for all $x,y\in[L_{\gamma s}]_{\gamma}$, all $n\in\mathbb{N}$ and $z\in[D^{n}_{\gamma}]_{\gamma}$.
Recall that $S_{\infty}$ is the set of states in $S_{\beta}$ for which $L_{\beta,s}]_{\beta}$ is infinite. If $s\in S_{\infty}$, then by Lemma  15 there exists a state $t=t(s)\in S_{\alpha}$ for which we may choose three non-collinear Gaussian integers such that each has a $(\alpha,D_{\alpha})$-expansion arriving at state $t$ and a $(\beta,D_{\beta})$-expansion arriving at state $s$, i.e., $x_{s,t},y_{s,t},z_{s,t}\in[L_{\beta,s}]_{\beta}\cap[L_{\alpha,t}]_{\alpha}$. Let
$\xi=\max\{\lvert x_{st}\rvert,\lvert y_{st}\rvert,\lvert z_{s,t}\rvert\>:\>s\in S_{\infty}\}+1,$
We take $\varepsilon=\frac{1}{\xi K}$ in Corollary 4, finding $m,n$ so that
(6) $\xi\lvert\alpha^{m}-\beta^{n}\rvert<\frac{1}{K}\lvert\beta^{n}\rvert$,
where $K$ is a natural number we will be free to choose when it is relevant. We shall fix these values of $m$ and $n$ from now on. An application of the triangle inequality gives
(7) $\lvert\alpha^{m}\rvert\geq\left(1-\frac{1}{\xi K}\right)\lvert\beta^{n}\rvert\geq\left(1-\frac{1}{K}\right)\lvert\beta^{n}\rvert$.
With $x_{s,t},y_{s,t},z_{s,t}$, we may also define two $\mathbb{R}$-linearly independent elements of $\mathbb{Z}[\mathrm{i}]$ for each $s\in S_{\infty}$, as follows:
(8) $p_{s,t}=(y_{s,t}-x_{s,t})(\alpha^{m}-\beta^{n}),\quad q_{s,t}=(z_{s,t}-x_{s,t})(\alpha^{m}-\beta^{n}).$
We claim that $|p_{s,t}|$ and $|q_{s,t}|$ are small relative to $|\beta^{n}|$. In particular
$\lvert p_{s,t}\rvert$ $=\lvert y_{s,t}-x_{s,t}\rvert\cdot\lvert\alpha^{m}-\beta^{n}\rvert$
$\leq(\lvert y_{s,t}\rvert+\lvert z_{s,t}\rvert)\cdot\lvert\alpha^{m}-\beta^{n}\rvert$
$\leq 2\xi\lvert\alpha^{m}-\beta^{n}\rvert$
$\leq\frac{2}{K}\lvert\beta^{n}\rvert$;
similarly for $\lvert q_{s,t}\rvert$.
We want to show that there exist a collection of balls $B$ for which $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ exhibits a given period $(p_{B},q_{B})$ in each ball $B$, and such that the balls cover most of $\mathbb{Z}[\mathrm{i}]$.
  • Claim. $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has step-period $(p_{s,t},q_{s,t})$ on the ball $B((x+1)\beta^{n},(1-\frac{2}{K})\beta^{n})$ for any $x\in[L_{\beta,s}]_{\beta}$.
[…]
    $\leq\lvert z-\beta^{n}\rvert+(\lvert y_{s,t}\rvert+1)\lvert\alpha^{m}-\beta^{n}\rvert$
    $\leq\left(1-\frac{2}{K}\right)\lvert\beta^{n}\rvert+\xi\cdot\frac{1}{\xi K}\lvert\beta^{n}\rvert$
    $\leq\left(1-\frac{1}{K}\right)\lvert\beta^{n}\rvert$
    $\leq\lvert\alpha^{m}\rvert$,
    where the last inequality follows from Equation (7).
    Now, since $x\in[L_{\beta,s}]_{\beta}$, this number must have at least one $(\beta,D_{\beta})$-expansion that ends at state $s$, so we can apply Lemma 12, and get the following:
    $a_{x\beta^{n}+z}$ $=a_{y_{s,t}\beta^{n}+z}$ $\text{since }x,y_{s,t}\in[L_{\beta,s}]_{\beta}\text{ and}$
[…]
    $=a_{y_{s,t}\alpha^{m}+z-y_{s,t}(\alpha^{m}-\beta^{n})}$ by adding and substracting $y_{s,t}\alpha^{m}$,
    $=a_{x_{s,t}\alpha^{m}+z-y_{s,t}(\alpha^{m}-\beta^{n})}$ $\text{since }x_{s,t},y_{s,t}\in[L_{\alpha,t}]_{\alpha}\text{ and}$
[…]
    $=a_{x_{s,t}\beta^{n}+z+p_{s,t}}$ by adding and substracting $x_{s,t}\beta^{n}$ and
    replacing $(x_{s,t}-y_{s,t})(\alpha^{m}-\beta^{n})$ by $p_{s,t}$,
    $=a_{x\beta^{n}+z+p_{s,t}}$ since $x_{s,t},x\in[L_{\beta,s}]_{\beta}$ and
    $\text{$z+p_{s,t}$ has a $(\beta,D_{\beta})$-expansion of length${}\leq n$},$
[…]
Therefore, by Lemma 19, $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has period $(p_{s,t},q_{s,t})$ on the ball $B((x+1)\beta^{n},(1-\frac{6}{K})\beta^{n})$ for any $x\in[L_{\beta,s}]_{\beta}$.
[…]
For any two balls $C_{u}$ and $B_{u+z}$ with $z\in\{\pm 1,\pm\mathrm{i}\}$, the distance between their centres will be $\lvert\beta^{n}\rvert$. Thus, Lemma 17, with $r_{1}=(1-\frac{6}{K})\beta^{n}$ and $r_{2}=(1-\frac{8}{K})\beta^{n}$, guarantees that their intersection contains a ball of radius $\frac{1}{2}(2-\frac{14}{K}-1)\lvert\beta^{n}\rvert=\frac{K-14}{2K}\lvert\beta^{n}\rvert$ .
As $|p_{u}|,|q_{u}|,|p_{u+z}|,|q_{u+z}|$ are each bounded by $\frac{2}{K}\lvert\beta^{n}\rvert$, if neither $u$ nor $u+z$ belong to $F$ we can apply Lemma 18, taking $U=C_{n}$ and $V=B_{n+z}$, if the intersection of these balls contains a ball of radius at least $\frac{4}{K}\lvert\beta^{n}\rvert$. Hence, we have the inequality:
$\frac{K-14}{2K}\lvert\beta^{n}\rvert\geq\frac{4}{K}\lvert\beta^{n}\rvert\implies K\geq 22,$
so any sufficiently large choice of $K$ allows us to apply this result and to conclude that $(a_{z})$ has a step-period $(p_{u},q_{u})$ in $C_{u}\cup B_{u+z}^{\circ(2\lvert\beta^{n}\rvert/K)}=C_{u}\cup C_{u+z}$.
Lemma 20 ensures that, whenever $K\geq 28$, the balls $\{C_{u}\>:\>u\not\in F\}$ cover the whole of $\mathbb{Z}[\mathrm{i}]$ except for some subset of $\left(\bigcup_{u\in F}C_{u}\right)\cap\mathbb{Z}[\mathrm{i}]$, which is a finite set, as it is a bounded subset of $\mathbb{Z}[\mathrm{i}]$. By enlarging the set $F$ if necessary, we may ensure that the induced subgraph of the Cayley graph of $(\mathbb{Z}[\mathrm{i}],+)$ with generating set $\{\pm 1,\pm\mathrm{i}\}$ is connected.
Fix any $u^{*}\in\mathbb{Z}[\mathrm{i}]\setminus F$, and let $v\in\mathbb{Z}[\mathrm{i}]\setminus F$ be any other element of this set. Let $u^{*}=u_{0},u_{1},\dotsc,u_{n}=v$ be a path between $u^{*}$ and $v$; since we see a step-period $(p_{u^{*}},q_{u^{*}})$ in $C_{u^{*}}$ and a period $(p_{u_{1}},q_{u_{1}})$ in $B_{u_{1}}$, an application of Lemma 18 shows that we observe a step-period $(p_{u^{*}},q_{u^{*}})$ in $C_{u_{1}}$; proceeding inductively, we show that a step-period $(p_{u^{*}},q_{u^{*}})$ in $C_{u_{j}}$ induces the same step-period on $C_{u_{j+1}}$. This applies, in particular, to $C_{v}$. As this is true for any $v\in\mathbb{Z}[\mathrm{i}]\setminus F$, we conclude that $(a_{z})_{z\in\mathbb{Z}[\mathrm{i}]}$ has a step-period $(p_{u^{*}},q_{u^{*}})$ in $\left(\bigcup_{u\in\mathbb{Z}[\mathrm{i}]\setminus F}C_{u}\right)\cap\mathbb{Z}[\mathrm{i}]$, which, as discussed before, equals the whole of $\mathbb{Z}[\mathrm{i}]$ bar a finite set. The result follows. ∎
As a corollary, we answer a question of [5].
Corollary 23.
Suppose that $U\subset\mathbb{Z}[\mathrm{i}]$ is a $\beta$-automatic set, where $|\beta|>1$. If $\overline{U}=U$, and $U$ is not eventually periodic, then $\beta$ is a root of an integer.
Proof.
Suppose that $U$ is $(\beta,D)$-automatic. If needed, using Proposition 10, we can enlarge $D$ so that it is closed under conjugation. Since $U=\overline{U}$, we conclude that $U$ is $(\bar{\beta},D)$-automatic. Since $U$ is not eventually periodic, $\beta$ and $\bar{\beta}$ are multiplicatively dependent. Thus there exist $m,n$ such that $|\beta|^{n}=|\bar{\beta}|^{m}$. This implies that we can take $m=n$, since $|\beta|>1$. Thus $\beta^{n}=\bar{\beta}^{n}$, so that $\beta^{n}\in\mathbb{R}$. ∎
Hansel and Safer were only interested in “natural” integral Gaussian numerations $(\beta,D)$, i.e., those where $D\subset\mathbb{N}$. The fact that $D$ is an integral digit set implies that $\beta\in\{-n\pm i:n\geq 1\}$ by Part (4) of Theorem 2. Thus given two multiplicatively independent $\alpha,\beta$ with integral digit sets, at least one of them is not a root of an integer. The following result now follows immediately from Theorem 1.
Corollary 24.
Let $\alpha$ and $\beta$ be two multiplicatively independent Gaussian integers with $|\alpha|,|\beta|>1$. Suppose that there exist digit sets $D_{\alpha}$ and $D_{\beta}$ consisting of natural numbers such that $(\alpha,D_{\alpha})$ and $(\beta,D_{\beta})$ are integral numeration systems. Then a configuration is $\alpha$- and $\beta$-automatic if and only if it is eventually periodic.
Finally, we mention the work of Cabezas and Petite [8] on constant shape substitutions. Via the representation of $\mathbb{Z}[\mathrm{i}]$ with the ring of integer matrices of the form $\begin{pmatrix}a&-b\\ b&a\end{pmatrix}$, a $\beta$-automatic configuration is also the coding of a fixed point of a constant shape substitution, determined by an integer $2\times 2$ matrix $M$ and fundamental domain $D$ for $M(\mathbb{Z}^{2})$. In the case where $M$ represents multiplication by $\beta\in\mathbb{Z}[\mathrm{i}]$, the configurations that Cabezas and Petite generate are $(\beta,D)$-automatic configurations. Our work therefore applies to their configurations, and in particular it means that if $\beta=a+ib$ is not the root of an integer, then their nontrivial $(\begin{pmatrix}a&-b\\ b&a\end{pmatrix},D)$-configurations cannot be replicated by $(k,k)$-automatic configurations, for $k\in\mathbb{N}$.
In conclusion, it is natural to wonder what happens for other number rings. A natural candidate is $\mathbb{Z}[\zeta]$, $\zeta$ a root of unity, or even quadratic integer rings.